[Halla_running] [Revised Logentry] Experiment status / summary of past 24 hours
efuchey at jlab.org
efuchey at jlab.org
Sun Apr 22 10:05:01 EDT 2018
Logentry Text:
--
Data taking continued smoothly over past 24 hours.
We had noticed in the logs some beam position drifts during April 21 Owl. Those are usually due to change configurations in Hall C which impact us by ricochet.
After that, shift crews were briefed to pay attention to this, and further such incidents were rectified quickly.
Yesterday evening, PS7 was also set to 0, to collect more interesting statistics. The interesting coincidence information is still present, albeit sunk into a lot of, ahem, random stuff.
Apparently, the T7 timing, introduced in the trigger, was messing the coincidence spectrum, and requires fine tuning (i.e. delay).
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3564975
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halla_running/attachments/20180422/72d811b5/attachment.html>
More information about the Halla_running
mailing list